Minimagic:

2. Over a period of 112 seconds, a wave crashes on shore 14 times. Between the crests of each wave is 3.2 x 1022 m. Use this information to determine the wave’s velocity.

Minimagic:

i need help bc i forgot how to find the numbers the correct way

Arieonna:

ok i think you need 112 seconds * 14 =1568 3.2*1022=3270.4 i hope this helps you

Minimagic:

@arieonna wrote:
ok i think you need 112 seconds * 14 =1568 3.2*1022=3270.4 i hope this helps you
but wont you have to divide 112 by 14?

Arieonna:

i don't think so if you see the word times behind a number i normally means multiply

Minimagic:

oh

Minimagic:

@arieonna wrote:
i don't think so if you see the word times behind a number i normally means multiply
what would the answer be? im confused bc there are two things up there.

Arieonna:

did you try to multiply 3270.4*1568=4838.4

Minimagic:

im sooooo confused now lmao one sec

Arieonna:

you can use a calculator to find what it multiplys to

Minimagic:

ohhhhhhhhhhhh wait its 3.2 x 10 to the 22 power

Minimagic:

@arieonna wrote:
you can use a calculator to find what it multiplys to
so its diff

Arieonna:

oh

Minimagic:

@arieonna wrote:
oh
can you still help? lol

TheKaylaHope:

@minimagic wrote:
i need help bc i forgot how to find the numbers the correct way
the wave's velocity is 4 x 10^21 m/s. Note that this is an extremely high velocity and is not physically possible for a wave in our current understanding of physics.